OpenStudy (anonymous):

let F(x)=1/t for t>0 Find the average rate of change of f over the interval (2,3) then repeat for the interval (2,T), where 2

zepdrix (zepdrix):

Hey there, \(\Large\rm f(t)=\dfrac{1}{t}\) Average rate of change from 2 to 3: \(\Large\rm \dfrac{f(3)-f(2)}{3-2}\)

zepdrix (zepdrix):

\(\Large\rm f(\color{royalblue}{t})=\dfrac{1}{\color{royalblue}{t}}\) \(\Large\rm f(\color{royalblue}{3})=\dfrac{1}{\color{royalblue}{3}}\)

zepdrix (zepdrix):

What'dyou think jp? :O Confusing stuff?

OpenStudy (anonymous):

the first part isnt too confusing then second part is mostly where im struggling with

OpenStudy (anonymous):

and thanks for the help

zepdrix (zepdrix):

Remember your slope formula? (y2-y1)/(x2-x1) This is the same idea, but now you're using function notation instead of y's. \(\Large\rm \dfrac{f(3)-f(2)}{3-2}=\dfrac{\frac{1}{3}-f(2)}{3-2}\) So we're evaluating the function at t=3, That's what f(3) is. And then we're plugging that value into our slope formula. How bout f(2)? What do you get when you plug 2 in for your t? (from the original function).

OpenStudy (anonymous):

if i plugged in 2 for my t ill get 1/2, so you saying after i get answer i have to put it into the slope formula?

zepdrix (zepdrix):

Mmmm, good. Yes. Plug it in. \(\Large\rm \dfrac{\frac{1}{3}-\frac{1}{2}}{3-2}\) Do you understand how to simplify this?

OpenStudy (anonymous):

find a common denominator so -1/6 divided by 1